in a cube close packed structure of mixed oxides ,the lattice is made up of oxide ions,one eighth of tetrahedral voids are occupied by divalent (X2+) ions,whileone half of the octahedral voids are occupied by trivalent ions (Y3+),the formula of oxide,?

Grade:12

5 Answers

Shriya Mehrotra
9 years ago
sir plz solve dis
Havishma Haranath
9 years ago
Actually the bigger molecules like oxygen occupy the lattice points.In ccp unit cell the effective number of atoms occupying lattice points is 8(1/8)+6(1/2)=4 and in ccp the number of tetrahedral voids is 8 so divalents ions will be 8/8 and number of octahedral voids are 4 so one half would be 2 finally,XY2O4.
britto george
9 years ago
this is easy question and you can try to solve it
Shruti
5 years ago
It can also be done by the use of spinal and inverse spinal structures..General formula is AB2O4 . where A is for divalent which will occupy tetrahedral voids , B for trivalent which will occupy octahedral voidsand O for CCP or FCC ..
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
The bigger molecules like oxygen occupy the lattice points.In ccp unit cell the effective number of atoms occupying lattice points = 8(1/8)+6(1/2) = 4
and in ccp the number of tetrahedral voids = 8
so divalents ions will be 8/8 and number of octahedral voids are 4 so its one half would be 2
finally, The formula is XY2O4.
 
Hope it helps.
Thanks and regards,
Kushagra